5.     

A car starts from rest and accelerates

uniformly over a time of 5.59 seconds for a distance of 150.00 m. Determine the

acceleration of the car.



 



6.     

Rocket-powered sleds are used to test

the human response to acceleration. If a rocket-powered sled is accelerated to

a speed of 360.00 m/s in 2.50 seconds, then what is the acceleration of the

rocket? What is the distance that the sled travels?


1
Expert's answer
2021-10-14T18:38:55-0400

5. Using the kinematic equation, obtain:


"d = \dfrac{at^2}{2}"

where "d = 150.00m" is the distance, "t = 5.59s" is the time, and "a" is the acceleration. Expressing "a", obtain:


"a = \dfrac{2d}{t^2} = \dfrac{2\cdot 150m}{(5.59s)^2} \approx 9.60m/s^2"

6. By definition of acceleration:


"a = \dfrac{v-v_0}{t}"

where "v = 360.00m\/s" is the final speed, "v_0 = 0" is the initial speed, and "t=2.50s" is the time. Thus, obtain:


"a = \dfrac{360m/s-0}{2.50s} = 144m/s^2"

The distance is determined with the same equation as in 5):


"d = \dfrac{at^2}{2} = \dfrac{144m/s^2\cdot (2.5s)^2}{2} = 450m"

Answer. 5) 9.60m/s^2, 6) 144m/s^2 and 450m.